The area of a circle is 64Pi. The area of a square is 256. Which is larger? the diameter of the circle or the side of the square?
Answer by checkley75(3666)

AREA=piR^2 OR
AREA=64pi OR
AREA=8pi WHERE THRE RADIUS ID 8 THUS THE DIAMETER IS 2*8=16
AREA OF A SQUARE IS X^2
X^2=256
X=16 WHICH THE LENGTH OF THE SIDE.
THEREFORE THE SIDE OF THE SQUARE (16) IS EQUAL TO THE DIAMETER (16) OF THE CIRCLE.